Which of the ordered pairs
(–4, 2), (4, 0), (–8, –3), (–16, 5)
are solutions for the equation x + 4y = 4?

Answer by Edwin McCravy(20055)



Let's see if (-4, 2) is a solution for 

x + 4y = 4

The first number in (-4, 2) is -4.  That tells you what to substitute for x.
The second number in (-4, 2) is 2.  That tells you what to substitute for y.

So in place of x in x + 4y = 4, write (-4) instead, and
in place of y in x + 4y = 4, write (2) instead.  Then we have this:

     x + 4y = 4
(-4) + 4(2) = 4
     -4 + 8 = 4
          4 = 4

Since 4 does equal 4 (-4, 2) IS a solution to x + 4y = 4.

-------------------
Let's see if (4, 0) is a solution for 

x + 4y = 4

The first number in (4, 0) is 4.  That tells you what to substitute for x.
The second number in (4, 0) is 0.  That tells you what to substitute for y.

So in place of x in x + 4y = 4, write (4) instead, and
in place of y in x + 4y = 4, write (0) instead.  Then we have this:

     x + 4y = 4
 (4) + 4(0) = 4
      4 + 0 = 4
          4 = 4

Since 4 does equal 4 (4, 0) IS a solution to x + 4y = 4.

-------------------
Let's see if (-8, -3) is a solution for 

x + 4y = 4

The first number in (-8, -3) is -8.  That tells you what to substitute for x.
The second number in (-8, -3) is -3.  That tells you what to substitute for y.

So in place of x in x + 4y = 4, write (-8) instead, and
in place of y in x + 4y = 4, write (-3) instead.  Then we have this:

      x + 4y = 4
(-8) + 4(-3) = 4
     -8 - 12 = 4
         -20 = 4

Since -20 does NOT equal 4 (-8, -3) is NOT a solution to x + 4y = 4.

-------------------

Let's see if (-16, 5) is a solution for 

x + 4y = 4

The first number in (-16, 5) is -16.  That tells you what to substitute for x.
The second number in (-16, 5) is 5.  That tells you what to substitute for y.

So in place of x in x + 4y = 4, write (-16) instead, and
in place of y in x + 4y = 4, write (5) instead.  Then we have this:

      x + 4y = 4
(-16) + 4(5) = 4
    -16 + 20 = 4
           4 = 4

Since 4 does equal 4 (-16, 5) IS a solution to x + 4y = 4.

------

So only one of them (-8, -3) is not a solution to x + 4y = 4.

Edwin